OpenStudy (anonymous):

find dy/dx given y=definite integral cos(e^t)dt. upperx^2, lower2. show steps/ explain

OpenStudy (loser66):

@SithsAndGiggles

OpenStudy (anonymous):

\[y=\int_2^{x^2}\cos\left(e^t\right)~dt\] In general, the fundamental theorem of calculus states that if \(f(x)=\displaystyle\int_{c}^{u(x)}f(t)~dt\), then its derivative is \(f'(x)=f(u(x))\cdot u'(x)\). (\(c\) is an arbitrary constant.) Here, \(f(t)=\cos\left(e^t\right)\) and \(u(x)=x^2\).

OpenStudy (anonymous):

thnx a million SithsAndGiggles :)

OpenStudy (anonymous):

You're welcome!

OpenStudy (loser66):

@SithsAndGiggles sorry for my dummy, I don't see the link between f(u(x) and f(t)

OpenStudy (anonymous):

ah no problem, thnx 4 at least trying

OpenStudy (loser66):

@unpi I am not the helper, I want to study from your problem, too.

OpenStudy (anonymous):

ok...

OpenStudy (loser66):

that's why I tag the person who I know he can help, not just help you, help me,too

OpenStudy (loser66):

ok, if you got it, show me, please

OpenStudy (anonymous):

k, give me a minute

OpenStudy (anonymous):

@Loser66, there's no elementary function that is equal to \[\int\cos\left(e^t\right)~dt\] But suppose there is, and let \(g(t)=\int\cos\left(e^t\right)~dt\). You can easily replace the \(t\) with \(x\) and it's the same function (the use of \(t\) in my other posts was as a dummy variable). By the FTC, the definite integral of the function we're working with looks like this, for some limits \(a\) and \(b\): \[\int_a^b\cos\left(e^t\right)~dt=g(b)-g(a)\] Let \(b=u(x)\). Then, \[\int_a^{u(x)}\cos\left(e^t\right)~dt=g(u(x))-g(a)\] Now take the derivative of both sides with respect to \(x\): \[\frac{d}{dx}\int_a^{u(x)}\cos\left(e^t\right)~dt=\frac{d}{dx}\left[g(u(x))-g(a)\right]\] Note that \(g(a)\) is a constant, so its derivative is 0: \[\frac{d}{dx}\int_a^{u(x)}\cos\left(e^t\right)~dt=\frac{d}{dx}g(u(x))\] By the chain rule, you have \[\frac{d}{dx}g(u(x))=g'(u(x))\cdot u'(x)\] Does that make sense?

OpenStudy (loser66):

yes, perfect, thanks a lot

OpenStudy (anonymous):

You're welcome